OpenStudy (anonymous):

Write an equation of a line that passes through the origin with slope 3. Please help!

OpenStudy (crashonce):

recall the equation (y-y1) = m(x-x1) sub in points

OpenStudy (crashonce):

y-0=3(x-0) therefore y=3x

OpenStudy (anonymous):

Thanks so much!

OpenStudy (godlovesme):

@cflores8613 Can u please close the question, if you are done getting help? Thanks! :)